OpenStudy (anonymous):

A vehical travels on a highway at a rate of 65 mi/h how long does it take the vehical to travel 25 mi? how do i set this up?

OpenStudy (luffingsails):

use the velocity formula and solve for time. \[v = \frac{d}{t}\] \[t = \frac{d}{v}\] Substitute your variables and solve.

OpenStudy (anonymous):

how bout this: 65 mi/hr = 65 / 3600 mi/second 25 mi / ( 65/3600) = 1384.6 seconds

OpenStudy (stormfire1):

Wow...there's a simpler way and you don't need to know any physics to do it: Simply divide 60 by 65 to get miles per minute...then multiply that by 25. You'll get ~23.08 minutes

OpenStudy (luffingsails):

There's no physics in applying a math formula. <chuckle> And you did exactly what the formula would have a person do.

OpenStudy (stormfire1):

That's true enough...I guess my point was that it's more of a ratio problem and doesn't require you to know any formulas.